= COS Problem 1: Recall that the Chebyshev nodes x4, x1,...,xy are determined on the interval (-1,1] as the zeros of Tn+1(x)class: numerical analysis

I wish if it was written in block letter

Sorry I can't read cursive
Answer #1

Answer:- Given that The interval [1,1] as the zeros of Tnt, (x) = cos(n+1) arcos () (2jt IT j=0,i,...n x; COS ht.12 f(x): Onm=1:- 24 T> (x) = cos (2cos () cos (06) - O coso - Cos(20) = 2 cos 0-1 2 2x = 1 T₂ (*) = 2X-1 from T2(x) = A (-4) (x-x*) 1 ja- cost 51 6 cos ( = cos(o-ho) COS 5 T3 (3) = cos (3005(x)) Cos00)=0 COSOX COS (30) = y coso - 3cOSE 433-34 42?_3x = (2-3) (3Rati santi < 2n 1 FC9) - Pr.Cu)) == CORD! Rnti 5nti an < R9+1 5nti 2n Inti (x) (n+1)! [by part o and Ifa (U) Isen] Cos (1) co